# Once a Triple, Not Always a Triple: The Evolution of Hierarchical Triples that Yield Merged Inner Binaries

### v1

@@ Abstract
More than half of all main-sequence (MS) stars have one or more companions, and many of those with initial masses M⊙ are born in hierarchical triples. These systems feature two stars in a close orbit (the inner binary) while a tertiary star orbits them on a wider orbit (the outer binary). In hierarchical triples, three-body dynamics combined with stellar evolution drives interactions and, in many cases, merges the inner binary entirely to create a renovated ‘Post-Merger Binary’ (PMB). By leveraging dynamical simulations and tracking binary interactions, we explore the outcomes of merged triples and investigate whether PMBs preserve signatures of their three-body history. Our findings indicate that in - of wide double WD binaries ( au), the more massive white dwarf (WD) is a merger product, implying that these DWD binaries were previously triples. Overall, we estimate that of observed wide DWDs originated in triple star systems and thereby have rich dynamical histories. Additionally, our results suggest that the separations of inner and outer binaries are uncorrelated at birth, providing insights into stellar formation processes. We also examine MS+MS and MS+Red Giant mergers manifesting as Blue Straggler stars (BSSs). These PMBs have orbital configurations and ages similar to most observed BSS binaries. While the triple+merger formation channel can explain the observed chemical abundances, moderate eccentricities, and companion masses in BSS binaries, it likely only accounts for - of BSSs. Meanwhile, we predict that the majority of observed single BSSs formed as collisions in triples and harbor long-period ( yr) companions. Furthermore, both BSS binaries and DWDs exhibit signatures of WD birth kicks.
@@ 1 Introduction
Observational and theoretical studies clearly show that most main-sequence (MS) stars form and evolve with one or more stellar companion (e.g., Raghavan et al., 2010; Tokovinin, 2014a; Moe & Di Stefano, 2017; Offner et al., 2023, see latter for a review). There are numerous possible processes that govern the early stages of formation. Fragmentation models suggest that most multiple star systems form via the gravitational collapse of over-dense regions that arise in cores, filaments, or massive accretion disks (e.g., Larson, 1972; Adams et al., 1989; Inutsuka & Miyama, 1992; Goodman et al., 1993; Kratter et al., 2008; Krumholz et al., 2009; Meyer et al., 2018). Dynamical capture, evolution, and migration models supply even more channels for the formation of multiple systems (e.g., Ostriker, 1994; Bate & Bonnell, 1997; Ostriker et al., 1999; Tohline, 2002; Moeckel & Bally, 2007; Bate, 2012; Lee et al., 2019; Cournoyer-Cloutier et al., 2021).
@@
During the evolution of hierarchical triples, the distant tertiary star will exchange angular momentum with the inner binary in a secular fashion, causing eccentricity-inclination oscillations: a process known as the eccentric Kozai-Lidov (EKL) mechanism (Kozai, 1962; Lidov, 1962; Naoz, 2016, see the latter for a review). Large eccentricities induced in the inner binary from EKL cycles lead triples to interact (i.e., transfer mass) far more often than they would as isolated binaries (e.g., Naoz & Fabrycky, 2014; Stephan et al., 2016; Cheng et al., 2019; Toonen et al., 2020; Shariat et al., 2023)111Other hierarchical perturbations, such as flyby stars, can also excite high eccentricities in wide systems, leading to interactions (e.g., Kaib & Raymond, 2014; Michaely & Perets, 2016, 2019, 2020; Michaely, 2021). This makes triples an especially interesting channel to study binary exotica such as Blue Straggler stars (BSSs; Geller et al., 2013), cataclysmic variables (CVs; Nelemans et al., 2001; Knigge et al., 2011; Pala et al., 2017; Knigge et al., 2022), and Post-Common Envelope Binaries (e.g., Toonen & Nelemans, 2013; Zorotovic et al., 2014; Hernandez et al., 2022; Yamaguchi et al., 2024a).
Beyond just mass transfer, interactions in the inner binary catalyzed by a distant companion can cause the inner binary to merge entirely (e.g., see Shariat et al., 2023, for solar-type stars). In such cases, the triple will become a binary where one star is a merger remnant and the other is the tertiary from the previous triple. We refer to these triples-turned-binaries as Post-Merger Binaries (PMBs).
@@ 2.1 Physical Processes and Numerical Setup
The initial semi-major axis distribution for Model IB is displayed in the top panel of Figure 1. We also show the initial (blue) and (black) distribution for the triples that led to a merged inner binary. The and distributions for Model IB closely resemble two Gaussians, whereas in the Model OB distribution is effectively Gaussian, while the distribution is a left-biased Gaussian constrained by the value. Overall, most inner binaries with close initial separations ( au) experienced a merger. Furthermore, the inner binaries with closer tertiaries, which would induce stronger EKL effects, were more likely to merge. In Figure 1, we also show the total number of triples (N) for each model at , the IMF information, and the eccentricity distribution (assumed uniform for both). Initial inclinations of the inner and outer binary were chosen from a uniform distribution.
@@ 3.2.1 Cooling Ages and Masses
To calculate the total age of an observed WD, Heintz et al. (2022) summed the best-fit progenitor age (its main-sequence) and the best-fit WD age based on its cooling track. Here, we calculate the best-fit progenitor age, they mapped the current mass of the WD to a progenitor’s ZAMS mass using a shifted MESA IFMR (Fields et al., 2016). From the ZAMS mass, one can then calculate the total main sequence lifetime, assuming single stellar evolution. If there were a merger during the star’s evolution, this age calculation method would give the wrong answer, leading to a discrepancy amongst the ages of stars in the binary DWDs. Since the study focuses on wide systems ( au, Heintz et al., 2022), the probability of previous mass transfer between the two stars is unlikely, making the merger channel a potentially robust explanation for the observed discrepancies. Here, we seek to understand whether the PMBs that became DWDs in our theoretical simulations also possess similarly discrepant cooling ages and therefore, support the merger channel.
@@
The right plot of Figure 3 shows the distribution of the primary WD mass in the DWD, normalized so that the area under the curve is unity. We display all observed DWDs in gray, the discrepant ones in red, the simulated post-merger DWDs in blue, and the discrepant simulated DWDs in dashed blue. Compared to all DWDs, the simulated post-merger DWDs have more massive primary WDs, which is similar to the discrepant sample. The observed discrepant sample sharply peaks at M⊙, similar to spectroscopic WD mass catalogs (e.g., Liebert et al., 2005; Tremblay et al., 2016). Our models peak at a similar mass and also contain a peak at slightly larger WD masses (- M⊙). The sharp maximum around M⊙ a function of the WD IFMR and the IMF. In particular, most IFMRs predict that MS progenitors with masses between M⊙ become M⊙ WDs (Fields et al., 2016), and most IMFs predict that a large fraction of MS stars fall in this range. This may indicate that observed WDs in this mass range originate from single stellar evolution or from a pre-WD merger that produced a main-sequence star in the M⊙ mass range (e.g., Liebert et al., 2005). The observed discrepant systems show a smaller peak around M⊙, which is consistent with being merger products (e.g., Liebert et al., 2005; Hallakoun et al., 2023). In the simulations, we find an additional local maximum at M⊙, which is not a feature that is observed with . This feature may be missing from observations because older and cooler WDs are too faint to be seen with . This high-mass peak may also be an artifact of the COSMIC models, and not be a physical manifestation of mergers (see Hansen et al., 2006; Tremblay et al., 2016, for example). Also, note that the simulated post-merger masses are heavily reliant on the post-merger mass calculation from COSMIC and our chosen Kroupa IMF.
@@ 4.2.2 Origin of the Orbital Structure of BSS Binaries
Already at , the three-body merger channel can successfully reproduce the orbital architectures of most observed BSS binaries. Observed systems with periods above days have periods and eccentricity that bode well with having previously been hierarchical triples. At this period range, the observed BSS binaries in NGC 188 (Geller et al., 2009) have a roughly uniform eccentricity distribution, similar to what is expected from the merger channel. On the other hand, the combined observations have a slight preference for smaller eccentricities () at this period range. These systems are closely similar to the RGMS mergers with similar periods of days but also are consistent with being post-Common Envelope binaries at this period range (Yamaguchi et al., 2024a). The preference for shorter eccentricities at these periods may also be a marker for a mass-transfer history within an isolated binary. Among the observed BSS binaries with days, most have non-zero, moderate eccentricities. This is surprising considering that they would have presumably circularized rapidly during their mass transfer evolution. A similar phenomenon is observed in post-Common Envelope binaries with similar periods (Yamaguchi et al., 2024a). For BSSs, this challenges the mass transfer formation channels and may indicate that another mechanism is pumping eccentricities. One eccentricity pumping mechanism to account for thetight BSS binaries with moderate eccentricities is WD formation kicks. When including the impact of WD kicks, we find that one of our short-period () BSS binaries with a WD companion escaped its circular orbit and became eccentric upon WD formation (bottom panel of Figure 5). Since the orbit is consistent with observations, kicks from WD companions can provide a method of explaining the eccentricities of tight BSS binaries. If this is true, we would predict that BSS binaries with orbital periods less than the circularization period and with non-zero eccentricities have WD companions. This scenario is already supported by observations from calculations of secondary masses to BSSs (Geller & Mathieu, 2011) and observations of hot WDs nearby BSSs (e.g., Gosnell et al., 2014, 2019; Sun et al., 2021; Jadhav et al., 2021; Vaidya et al., 2022; Nine et al., 2023).
@@
From COSMIC, an MS+MS merger generally manifests as one inflated MS star, which would have highly efficient tides and an assumed zero eccentricity. Notably, some of the MSMS merger systems sparked a short period of mass transfer during their inflated state or later on as the tertiary evolved off the main sequence. The latter scenario is the same as shown on the right side of Figure 2. Such a scenario is an example of triple mass transfer, which has been noted previously (Gao et al., 2023; Nine et al., 2024), and we find here that it can be a fascinating channel to form close-in BSS binaries.
@@
Overall, we find that the triple+merger channel can reproduce the periods, eccentricities, and many other discussed characteristics of most observed BSS binaries. However, observations of BSS binaries clearly show that - of them have companions with orbital periods less than days (Geller & Mathieu, 2011; Geller et al., 2015). In the triple+merger channel, most tertiaries lie at periods above this range and beyond the RV detection limit of the BSS surveys, making them observable as single BSSs. Therefore, for BSS binaries with periods below days, the stable mass transfer channel is likely to be dominant. Nevertheless, there exists a handful of observed BSS binaries with peculiar properties in this period regime that can be naturally explained by the triple+merger channel or by accounting for additional binary evolution mechanisms. Furthermore, based on our models, it is likely that the - of BSSs without nearby companions formed via the triple+merger channel and have tertiary stars beyond the detection limit.
@@ 7 Conclusions
Formation channel for BSS binaries: The orbital configuration of PMBs that had MSMS or RGMS mergers closely matches with most observed BSS binaries (Figure 5). Based on this, we predict that the - of BSSs without nearby companions formed via stellar collisions and have very long period ( days) companions. However, current observations show that most observed BSSs in clusters have close companions ( days; Geller & Mathieu, 2011; Geller et al., 2015), which supports that a majority of BSSs formed via stable mass transfer. However, the eccentricities in some of the close BSS binaries are inconsistent with stable mass transfer, but can naturally be explained by triple dynamics or by considering an additional eccentricity pumping mechanism (e.g., WD kicks). We also show that collisions in hierarchical triples often occur after Myr to Gyr of triple evolution, producing BSS ages consistent with old clusters. Lastly, similar to the mass transfer channel, the triple+merger channel would also produce enriched chemical abundances, fast rotation rates, and companion masses observed in many BSSs.
@@
We thank Tyler M. Heintz and J.J. Hermes for bringing our attention to the age-discrepant DWDs in the Gaia sample. This work used computational and storage services associated with the Hoffman2 Shared Cluster provided by UCLA Office of Advanced Research Computing’s Research Technology Group. S.N. acknowledges the partial support from NASA ATP 80NSSC20K0505 and from the NSF-AST 2206428 grant, as well as thanks Howard and Astrid Preston for their generous support.

### v2

@@ Abstract
More than half of all main-sequence (MS) stars have one or more companions, and many of those with initial masses M⊙ are born in hierarchical triples. These systems feature two stars in a close orbit (the inner binary) while a tertiary star orbits them on a wider orbit (the outer binary). In hierarchical triples, three-body dynamics combined with stellar evolution drives interactions and, in many cases, merges the inner binary entirely to create a renovated ‘Post-Merger Binary’ (PMB). By leveraging dynamical simulations and tracking binary interactions, we explore the outcomes of merged triples and investigate whether PMBs preserve signatures of their three-body history. Our findings indicate that in - of wide double WD binaries ( au), the more massive white dwarf (WD) is a merger product, implying that these DWD binaries were previously triples. Overall, we estimate that of observed wide DWDs originated in triple star systems and thereby have rich dynamical histories. We also examine MS+MS and MS+Red Giant mergers manifesting as Blue Straggler stars (BSSs). These PMBs have orbital configurations and ages similar to most observed BSS binaries. While the triple+merger formation channel can explain the observed chemical abundances, moderate eccentricities, and companion masses in BSS binaries, it likely only accounts for - of BSSs. Meanwhile, we predict that the majority of observed single BSSs formed as collisions in triples and harbor long-period ( yr) companions. Furthermore, both BSS binaries and DWDs exhibit signatures of WD birth kicks.
@@ 1 Introduction
Observational and theoretical studies clearly show that most main-sequence (MS) stars form and evolve with one or more stellar companion (e.g., Raghavan et al., 2010; Tokovinin, 2014a; Moe & Di Stefano, 2017; Offner et al., 2023, see latter for a review). For multiple star systems, there are numerous possible processes that govern the early stages of formation. Fragmentation models suggest that most multiple star systems form via the gravitational collapse of over-dense regions that arise in cores, filaments, or massive accretion disks (e.g., Larson, 1972; Adams et al., 1989; Inutsuka & Miyama, 1992; Goodman et al., 1993; Kratter et al., 2008; Krumholz et al., 2009; Meyer et al., 2018). Dynamical capture, evolution, and migration models supply even more channels for the formation of multiple systems (e.g., Ostriker, 1994; Bate & Bonnell, 1997; Ostriker et al., 1999; Tohline, 2002; Moeckel & Bally, 2007; Bate, 2012; Lee et al., 2019; Cournoyer-Cloutier et al., 2021).
@@
During the evolution of hierarchical triples, the distant tertiary star will exchange angular momentum with the inner binary in a secular fashion, causing eccentricity-inclination oscillations: a process known as the eccentric Kozai-Lidov (EKL) mechanism (von Zeipel, 1910; Kozai, 1962; Lidov, 1962; Naoz, 2016, see the latter for a review). Large eccentricities induced in the inner binary from EKL cycles lead triples to interact (i.e., transfer mass) far more often than they would as isolated binaries (e.g., Naoz & Fabrycky, 2014; Stephan et al., 2016; Cheng et al., 2019; Toonen et al., 2020; Shariat et al., 2023)111Other hierarchical perturbations, such as flyby stars, can also excite high eccentricities in wide systems, leading to interactions (e.g., Kaib & Raymond, 2014; Michaely & Perets, 2016, 2019, 2020; Michaely, 2021). This makes triples an especially interesting channel to study binary exotica such as Blue Straggler stars (BSSs; Geller et al., 2013), cataclysmic variables (CVs; Nelemans et al., 2001; Knigge et al., 2011; Pala et al., 2017; Knigge et al., 2022), and Post-Common Envelope Binaries (e.g., Toonen & Nelemans, 2013; Zorotovic et al., 2014; Hernandez et al., 2022; Yamaguchi et al., 2024a).
Beyond just mass transfer, interactions in the inner binary catalyzed by a distant companion can cause the inner binary to merge entirely (e.g., Antonini et al., 2017; Grishin & Perets, 2022; Trani et al., 2022; Preece et al., 2022; Rajamuthukumar et al., 2023; Shariat et al., 2023; Dorozsmai et al., 2024). In such cases, the triple will become a binary where one star is a merger remnant and the other is the tertiary from the previous triple. We refer to these triples-turned-binaries as Post-Merger Binaries (PMBs).
@@ 2.1 Physical Processes and Numerical Setup
The initial semi-major axis distribution for Model IB is displayed in the top panel of Figure 1. We also show the initial (blue) and (black) distribution for the triples that led to a merged inner binary. The and distributions for Model IB closely resemble two Gaussians, whereas in the Model OB distribution is effectively Gaussian, while the distribution is a left-biased Gaussian constrained by the value. Overall, most inner binaries with close initial separations ( au) experienced a merger. Furthermore, the inner binaries with closer tertiaries, which would induce stronger EKL effects, were more likely to merge. In Figure 1, we also show the total number of triples (N) for each model at , the IMF information, and the eccentricity distribution (assumed uniform for both). The initial inner and outer inclinations are chosen from an isotropic distribution.
@@ 3.2.1 Cooling Ages and Masses
To calculate the total age of an observed WD, Heintz et al. (2022) summed the best-fit progenitor age (its main-sequence lifetime) and the best-fit WD age based on its cooling track. To calculate the best-fit progenitor age, they mapped the current mass of the WD to a progenitor’s ZAMS mass using a shifted MESA IFMR (Fields et al., 2016). From the ZAMS mass, one can then calculate the total main sequence lifetime, assuming single stellar evolution. If there were a merger during the star’s evolution, this age calculation method would give the wrong answer, leading to a discrepancy amongst the ages of stars in the binary DWDs. Since the study focuses on wide systems ( au, Heintz et al., 2022), the probability of previous mass transfer between the two stars is unlikely, making the merger channel a potentially robust explanation for the observed discrepancies. Here, we seek to understand whether the PMBs that became DWDs in our theoretical simulations also possess similarly discrepant cooling ages and therefore, support the merger channel.
@@
The right plot of Figure 3 shows the distribution of the primary WD mass in the DWD, normalized so that the area under the curve is unity. We display all observed DWDs in gray, the discrepant ones in red, the simulated post-merger DWDs in blue, and the discrepant simulated DWDs in dashed blue. Compared to all DWDs, the simulated post-merger DWDs have more massive primary WDs, which is similar to the discrepant sample. The observed discrepant sample sharply peaks at M⊙, similar to spectroscopic WD mass catalogs (e.g., Liebert et al., 2005; Tremblay et al., 2016). Our models peak at a similar mass and also contain a peak at slightly larger WD masses (- M⊙). The sharp maximum around M⊙ is an attribute of the WD IFMR and the IMF. Namely, most IFMRs predict that MS progenitors with masses between M⊙ become M⊙ WDs (Fields et al., 2016), and most MS stars that are now WDs had initial masses in this range. This may indicate that observed WDs in this mass range originate from single stellar evolution or from a pre-WD merger that produced a main-sequence star in the M⊙ mass range (e.g., Liebert et al., 2005). The observed discrepant systems show a smaller peak around M⊙, which is consistent with being merger products (e.g., Liebert et al., 2005; Hallakoun et al., 2023). In the simulations, we find an additional local maximum at M⊙, which is not a feature that is observed with . This feature may be missing from observations because older and cooler WDs are too faint to be seen with . This high-mass peak may also be an artifact of the COSMIC models, and not be a physical manifestation of mergers (see Hansen et al., 2006; Tremblay et al., 2016, for example). Also, note that the simulated post-merger masses are heavily reliant on the post-merger mass calculation from COSMIC and our chosen Kroupa IMF.
@@ 4.2.2 Origin of the Orbital Structure of BSS Binaries
At , the three-body merger channel can successfully reproduce the orbital architectures of most systems, except the shortest BSS binaries. Notably, observed systems with periods above days have periods and eccentricity that bode well with having previously been hierarchical triples. At this period range, the observed BSS binaries in NGC 188 (Geller et al., 2009) have a roughly uniform eccentricity distribution, similar to what is expected from the merger channel. On the other hand, the combined observations have a slight preference for smaller eccentricities () at this period range. These systems are closely similar to the RGMS mergers with similar periods of days but also are consistent with being post-Common Envelope binaries at this period range (Yamaguchi et al., 2024a). The preference for shorter eccentricities at these periods may also be a marker for a mass-transfer history within an isolated binary. Among the observed BSS binaries with days, most have non-zero, moderate eccentricities. This is surprising considering that they would have presumably circularized rapidly during their mass transfer evolution. A similar phenomenon is observed in post-Common Envelope binaries with similar periods (Yamaguchi et al., 2024a). For BSSs, this challenges the mass transfer formation channels and may indicate that another mechanism is pumping eccentricities.
One eccentricity pumping mechanism to account for the tight BSS binaries with moderate eccentricities is WD formation kicks. When including the impact of WD kicks, we find that one of our short-period () BSS binaries with a WD companion escaped its circular orbit and became eccentric upon WD formation (bottom panel of Figure 5). Since the orbit is consistent with observations, kicks from WD companions can provide a method of explaining the eccentricities of tight BSS binaries. If this is true, we would predict that BSS binaries with orbital periods less than the circularization period and with non-zero eccentricities have WD companions. This scenario is already supported by observations from calculations of secondary masses to BSSs (Geller & Mathieu, 2011) and observations of hot WDs nearby BSSs (e.g., Gosnell et al., 2014, 2019; Sun et al., 2021; Jadhav et al., 2021; Vaidya et al., 2022; Nine et al., 2023).
@@
From COSMIC, an MS+MS merger generally manifests as one inflated MS star, which would have highly efficient tides and an assumed zero eccentricity. Notably, some of the MSMS merger systems sparked a short period of mass transfer during their inflated state or later on as the tertiary evolved off the main sequence. The latter scenario is the same as shown on the right side of Figure 2. Such a scenario is an example of triple (or tertiary) mass transfer, which has been noted previously (Gao et al., 2023; Nine et al., 2024; Dorozsmai et al., 2024), and we find here that it can be a fascinating channel to form close-in BSS binaries.
@@
Overall, we find that the triple+merger channel can reproduce the periods, eccentricities, and many other discussed characteristics of most observed BSS binaries. However, observations of BSS binaries clearly show that - of them have companions with orbital periods less than days (Geller & Mathieu, 2011; Geller et al., 2015). In the triple+merger channel, most tertiaries lie at periods above this range and beyond the RV detection limit of the BSS surveys, making them observable as single BSSs. Therefore, for BSS binaries with periods below days, the stable mass transfer channel is likely to be dominant. Nevertheless, there exists a handful of observed BSS binaries with peculiar properties in this period regime that can be naturally explained by the triple+merger channel or by accounting for additional binary evolution mechanisms. Furthermore, based on our models, it is likely that the - of BSSs without nearby companions formed via the triple+merger channel and have tertiary stars beyond the detection limit. Previous studies have estimated that the fraction of BSSs formed in triples is closer to (Antonini et al., 2016). Lastly, it is important to note that mass transfer will not always lead to circular orbits, which is a feature of COSMIC models (e.g., Hamers & Dosopoulou, 2019).
@@ 7 Conclusions
Formation channel for BSS binaries: The orbital configuration of PMBs that had MSMS or RGMS mergers closely matches with most ( days) observed BSS binaries (Figure 5). Based on this, we predict that the - of BSSs without nearby companions formed via stellar collisions and have very long period ( days) companions. However, current observations show that most observed BSSs in clusters have close companions ( days; Geller & Mathieu, 2011; Geller et al., 2015), which supports that a majority of BSSs formed via stable mass transfer. Still, the eccentricities in some of the close BSS binaries are inconsistent with stable mass transfer, but can naturally be explained by triple dynamics or by considering an additional eccentricity pumping mechanism (e.g., WD kicks). We also show that collisions in hierarchical triples often occur after Myr to Gyr of triple evolution, producing BSS ages consistent with old clusters. Lastly, similar to the mass transfer channel, the triple+merger channel would also produce enriched chemical abundances, fast rotation rates, and companion masses observed in many BSSs.
@@
We thank Tyler M. Heintz and J.J. Hermes for bringing our attention to the age-discrepant DWDs in the Gaia sample. We thank the anonymous referee for a constructive report. This work used computational and storage services associated with the Hoffman2 Shared Cluster provided by UCLA Office of Advanced Research Computing’s Research Technology Group. S.N. acknowledges the partial support from NASA ATP 80NSSC20K0505 and from the NSF-AST 2206428 grant, as well as thanks Howard and Astrid Preston for their generous support.